A uniform magnetic field equal to 1.48 T k is in the +z direction. Find the force exerted by the field on the protein if the velocity of the protein is (a) 2.7 km/s i (b) 3.7 km/s j (c) 6.8 km/s k and (d) 4.0 km/s i + 3.0 km/s j.
Answer #1

a)

q = 1.602 * 10^-19 C

Fb = q ( v * B)

Fb = 1.602 * 10^-19 * 3.996 * 10^3

Fb = 6.40 * 10^-16 N

b)

Fb = 1.602 * 10^-19 *5.476 *10^3

Fb = 8.772 * 10^-16 N

c)

Fb = 0 N

d)

magnitude of ( v *B) = sqrt(-4.44^2 + 5.92^2) = 3.91

then

Fb = 1.602 * 10^-19 * 3.915 * 10^3

Fb = 6.27 * 10^-16 N